Evaluate the function f(x) = 4x + 6 for x= 4
Answer by nerdybill(7384)

Evaluate the function f(x) = 4x + 6 for x= 4
.
simply replace x with 4 and solve:
starting with:
f(x) = 4x + 6
substitute our value for x:
f(4) = 4(4) + 6
f(4) = 16 + 6
f(4) = 22